# Learning Fine Pinch-Grasp Skills using Tactile Sensing from Real Demonstration Data

### v1

@@ IV-D Results of Grasping Tasks
The BC network trained on human demonstration data is deployed on a real dual-arm robot to verify its performance by the designed tasks. A set of snapshots showing grasping a tube from the tube holder and the table is shown in Fig. 6a. In both grasping tasks, the learned control policy achieved $100\%$ success rate, even when the initial poses of the tube are different from their original pose in the demonstration.
The dual-arm robot can make prompt adjustments and enable stable dexterous grasping by learning from only one demonstration. In the process of lifting the object, the dual-arm robot achieves stable grasping by constantly twiddling the “fingertips” (tips of TacTip sensors) and adjusting the object to the central position. The process of retrieving an object from the table and adjusting its pose to maintain balance requires very fine movements and interactions supported by rich tactile information, where a 6-axis force/torque information is not sufficient to discern different contact situations in this scenario.
We evaluate the robustness of the learned policy against external disturbances. It can be seen from Fig. 4a that the dual-arm robot can make a proper adjustment to adapt to pushes. Although the pose of the two-arm robot in contact with the object was changed each time while being pushed, the dual-arm robot can always fine-adjust the object reactively to the center of the fingertips (Tactip sensors), roll and move the object to the desired position. Compared with the manually programmed behavior, this serves as a feedback policy that has been successfully acquired from human dexterity skills, which enables the dual-arm robot to autonomously adjust the posture and ensure a stable grasp quickly. It is noteworthy that such active rolling adjustment has not been explicitly demonstrated by any separate trials, but rather, this behavior was successfully captured by the rich tactile data during one-shot demonstration of pick-lift grasping.
To examine the reaction in the presence of an unknown situation, i.e., grasping failures, the learned policy demonstrated contract-awareness of the falling object, i.e., loss of contact according to the tactile feedback, and thus controls the robot to restart the grasping process, which was not explicitly programmed or demonstrated by the prior LfD data. The result of the regrasping experiments in Fig. 4b shows that the tactile-based control learned from human demonstrations is very effective in performing robotic dexterous bimanual manipulation tasks autonomously and quickly without the need for explicit manual programming or complex planning.
@@ IV-E Comparison Study
We conducted a comparison study to validate that successful grasping is achieved by the active use of tactile sensing. Besides training a BC network using the structure shown in Fig. 3b, we also train two different BC networks for comparison. The first one has exactly the same BC network structure but with frozen input of the tactile images, meaning that the image input stays unchanged during both the training and testing. The second one has an FCN structure and uses the poses of two end-effectors (both positions and orientation) as the input to train the network. The proposed BC network demonstrates convergence to a loss of 0.04 on the testing set. In contrast, the network employing frozen tactile information achieves convergence with a loss of 0.5, while the FCN converges to a loss value of 1. These results prove that the effective integration of tactile information significantly enhances the convergence rate and leads to a reduced loss value in the final model.
We also compared all the grasping performances on the real dual-arm robot. As shown in Fig. 5, both BC network structures without using the tactile information failed in grasping the tube: robot arms failed to approach the object from its initial pose, and instead, they bypassed the object and moved towards the desired end-poses, showing no contact-awareness. The experimental results indicate that tactile feedback plays an essential role in providing contact information for initiating contacts, generating appropriate adjustments, lifting and retrieving to the desired target locations, enabling the dual-arm robot to perform very fine and dexterous contact-rich skills.
@@ IV-F Interpretability
where $W_{i}$ and $W_{j}$ are the weight distributions of each modality. $N(\cdot)$ represents the normalization process. $I$ is importance of the tactile information that is calculated by adding all the absolute value of weight that the learned policy distributed to tactile features. $J$ is calculated in the same way by adding all the the absolute value of weight that are distributed to robot proprioceptive state features.
The comprehensive process of dexterous pinch grasping can be subdivided into four primary stages: pre-grasp, pressing, rolling and lifting, and stabilization. Each of these stages utilizes tactile feedback in a distinct manner. In Fig 6b, the weight changes during the complete dexterous pinch grasping process are depicted. Initially, as the end-effector moves toward the objects without any contact deformation on the tactile sensor, the weight of the robot’s proprioceptive state exceeds that of the tactile information. When the tactile sensor comes into contact with the desk and is prepared for a pre-grasp pose, the weight of the tactile information increases (stage A). As the end-effector advances towards the object and initiates contact, the weight attributed to the tactile information increase, exceeding that of the proprioceptive state (stage B). During the roll and lift phase, the weight of the tactile information initially decreases, subsequently achieving equilibrium with the proprioceptive state (stage C). This indicates that during the lifting phase, the learned policy necessitates tactile information for successful in-hand manipulation and proprioceptive information for effective dual-arm coordination. Finally, upon successfully lifting the tube, the weight reverts to the tactile information, facilitating the stabilization of the tube (stage D).
@@ V Conclusion And Future Work
The presented Learning from Demonstration (LfD) framework showed successful skill transfer from humans to robots with minimal one-trial of real robot data, with the use of rich tactile sensing at robot’s fingertips. Through our journey of exploring how to best utilize the new generation of compliant tactile sensors, we have developed the presented encoding methods that can effectively extract and capture high-dimensional contact sensing from soft tactile sensors, together with the fusion with proprioceptive feedback. The interesting outcome is to confirm the possibility to learn from real robot data directly, without the need of high computation or big data, if the right data is used.
Our comparison studies showed that without the use of tactile sensing, dexterous motor skills cannot be learned by one-shot demonstrations with traditional robot sensing which is rather limited. Our proposed approach overcomes the traditional limitations of one-shot learning method, through the use of tactile and proprioceptive information for extracting useful information and mapping it into fine-grained motor skills. This approach is shown to be robust in the presence of external pushes and is able to perform re-grasp the object if it drops, which was not shown in the one-shot demonstration and emerges as the natural outcome of sensorimotor skills through state-action mapping. The ability to learn from real data/hardware and a single demonstration is very attractive to gain a wider range of machine learning approaches in the real world, where tasks can hard be simulated and only a small amount of data is available.
Meanwhile, one apparent limitation is that one-shot learning is apriori trained on a specific task and object, and it can be generalised and robust only around neighbourhood situations within a category of similar tasks: generalization applies to new/unseen objects that are similar to the demonstrated object of certain variations. The advantage of having only one demonstration comes with the trade-off that when a very different object grasping is needed, then at least one demonstration data is needed. Another limitation is that the robot’s performance is based on blind grasping and re-grasping, and has not yet utilised external visual perception. In the future, integration of the current framework with stereo vision could extend the versatility and dexterity of object manipulation. Overall, our proposed LfD framework provides an attractive solution for learning from one demonstration with tactile sensing and supports broad real-world applications in robotics with data scarcity.

### v2

@@ IV-D Results of Grasping Tasks
The BC network trained on human demonstration data is deployed on a real dual-arm robot to verify its performance by the designed tasks. In both grasping tasks, the learned control policy achieved a $95\%$ success rate, even when the initial poses of the tube were different from their original pose in the demonstration. Occasionally, the policy might not succeed in grasping the object on the first attempt. However, it consistently adapts the behavior based on the contact situation with objects and will attempt to repeat the grasping action upon failure. The policy fails to grasp an object only when the object falls outside the workspace boundaries.
The dual-arm robot can make prompt adjustments and enable stable dexterous grasping by learning from only few demonstrations. In the process of lifting the object, the dual-arm robot achieves stable grasping by constantly twiddling the “fingertips” (tips of TacTip sensors) and adjusting the object to the central position. The process of retrieving an object from the table and adjusting its pose to maintain balance requires very fine movements and interactions supported by rich tactile information, where a 6-axis force/torque information is not sufficient to discern different contact situations in this scenario.
We evaluate the robustness of the learned policy against external disturbances. It can be seen from Fig. 3a that the dual-arm robot can make a proper adjustment to adapt to pushes. Although the pose of the two-arm robot in contact with the object was changed each time while being pushed, the dual-arm robot can always fine-adjust the object reactively to the center of the fingertips (Tactip sensors), roll and move the object to the desired position. Compared with the manually programmed behavior, this serves as a feedback policy that has been successfully acquired from human dexterity skills, which enables the dual-arm robot to autonomously adjust the posture and ensure a stable grasp quickly. It is noteworthy that such active rolling adjustment has not been specially demonstrated by any separate trials, but rather, this behavior was successfully captured by the rich tactile data during the demonstration of pick-lift grasping.
To examine the reaction in the presence of an unknown situation, i.e., grasping failures, the learned policy demonstrated contract-awareness of the falling object, i.e., loss of contact according to the tactile feedback, and thus controls the robot to restart the grasping process, which was not explicitly programmed or demonstrated by the prior LfD data. The result of the re-grasping experiments in Fig. 3b shows that the tactile-based control learned from human demonstrations is very effective in performing robotic dexterous bimanual manipulation tasks autonomously and quickly without the need for explicit manual programming or complex planning.
@@ IV-E Comparison Study
We conducted a comparison study to validate that successful grasping is achieved by the active use of tactile sensing. Besides training a BC network using the structure shown in Fig. 2, we also train two different BC networks for comparison. The first one has exactly the same BC network structure but with frozen tactile images as input to CAE, meaning that the encoded image feature input stays unchanged during both the training and testing. The second one has an FCN structure and uses the poses of two end-effectors (both positions and orientation) as the input to train the network. The proposed BC network demonstrates convergence to a loss of 0.04 on the testing set. In contrast, the network employing frozen tactile information achieves convergence with a loss of 0.5, while the FCN converges to a loss value of 1. These results prove that the effective integration of tactile information significantly enhances the convergence rate and leads to a reduced loss value in the final model.
We also compared all the grasping performances of the real dual-arm robot. As shown in Fig. 4, both BC network structures without using the tactile information failed in grasping the tube: robot arms failed to approach the object from its initial pose, and instead, they bypassed the object and moved towards the desired end-poses, showing no contact-awareness. The experimental results indicate that tactile feedback plays an essential role in providing contact information for initiating contacts, generating appropriate adjustments, lifting and retrieving to the desired target locations, enabling the dual-arm robot to perform very fine and dexterous contact-rich skills.
@@ IV-F Interpretability
where $W_{i}$ and $W_{j}$ are the weight distributions of each modality. $N(\cdot)$ represents the normalization process. $I$ is the importance of the tactile information that is calculated by adding all the absolute values of weight that the learned policy distributed to tactile features. $J$ is calculated in the same way by adding all the absolute values of weight that are distributed to robot proprioceptive state features.
The comprehensive process of dexterous pinch grasping can be subdivided into four primary stages: pre-grasp, pressing, rolling and lifting, and stabilization. Each of these stages utilizes tactile feedback in a distinct manner. In Fig 5b, the weight changes during the complete dexterous pinch grasping process are depicted. Initially, as the end-effector moves toward the objects without any contact deformation on the tactile sensor, the weight of the robot’s proprioceptive state exceeds that of the tactile information. When the tactile sensor comes into contact with the desk and is prepared for a pre-grasp pose, the weight of the tactile information increases (stage A). As the end-effector move towards the object and initiates contact, the weight attributed to the tactile information increases, exceeding that of the proprioceptive state (stage B). During the roll and lift phase, the weight of the tactile information initially decreases, subsequently achieving equilibrium with the proprioceptive state (stage C). This indicates that during the lifting phase, the learned policy necessitates both tactile information for successful in-hand manipulation and proprioceptive information for effective dual-arm coordination. Finally, upon successfully lifting the tube, the weight reverts to the tactile information, facilitating the stabilization of the tube (stage D).
@@ V Conclusion And Future Work
In this work, we introduce a tactile-driven LfD framework that demonstrates promising results in bimanual pinch grasping with a limited number of real robot demonstrations. Our exploration into leveraging the latest compliant tactile sensors has led to the development of the presented encoding methods that can effectively extract and capture high-dimensional contact sensing from soft tactile sensors, together with the fusion with proprioceptive feedback. The interesting outcome is to confirm the possibility of learning from real robot data directly, eliminating the necessity for large datasets and extensive training time, if the right data is effectively used.
Our comparison studies showed that without the use of tactile sensing, dexterous motor skills cannot be learned by few-shot demonstrations with traditional robot sensing which is rather limited. In contrast, our approach demonstrates remarkable robustness in the presence of external pushes and is able to perform re-grasp the object if it drops. This ability was not explicitly illustrated in the initial demonstrations, emerging instead as a natural consequence of contact-aware sensorimotor skills through state-action mapping.
Meanwhile, one apparent limitation is that the skill needs to be trained on a specific task, and it can be generalised and robust only around neighbourhood situations within a category of similar tasks: generalization applies to new/unseen objects that are similar to the demonstrated object of certain variations. Another limitation is that the robot’s performance is based on blind grasping and re-grasping, and has not yet utilized external visual perception. In the future, integration of the current framework with stereo vision could extend the versatility and dexterity of object manipulation. Overall, our proposed LfD framework provides an attractive solution for learning from a few demonstrations with tactile sensing and supports broad real-world applications in contact-rich manipulation tasks.